# Linear Convergence of the Proximal Gradient Method for Composite Optimization Under the Polyak-Łojasiewicz Inequality and Its Variant

### v1

@@ Abstract
We study the linear convergence rates of the proximal gradient method for composite functions satisfying two classes of Polyak-Łojasiewicz (PL) inequality: the PL inequality, the variant of PL inequality defined by the proximal map-based residual. Using the performance estimation problem, we either provide new explicit linear convergence rates or improve existing complexity bounds for minimizing composite functions under the two classes of PL inequality.
@@ 1 Introduction
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is a function with Lipschitz continuous gradient, and $h:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is a closed, proper and convex function. In addition, we assume that the optimal set $X_{*}(F):=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}F(x)$ is nonempty. We denote the class of closed and proper convex functions by $\mathcal{F}_{0,\infty}$, the class of $L$-smooth convex functions by $\mathcal{F}_{0,L}$, and the class of $L$-smooth functions by $\mathcal{F}_{-L,L}$. For simplicity, we say that $F$ is a nonconvex composite function if $f\in\mathcal{F}_{-L,L}$ and a convex composite function if $f\in\mathcal{F}_{0,L}$.
@@
2:for $k=1,\ldots,N$ do
3:     $x_{k+1}=\mathrm{prox}_{\gamma h}\left(x_{k}-\gamma\nabla f(x_{k})\right)$
4:     $k=k+1$
5:end for
@@
It is well-known that for composite functions with strongly convex $f$, the PGM achieves a linear convergence rate [2]. Relaxing the strong convexity condition while retaining a linear convergence rate of the PGM has been extensively studied in the literature. The PL inequality was initially introduced by Polyak [13], which is used to show linear convergence of gradient descent. Bolte et al. [3] derived a linear convergence rate of the PGM under the Kurdyka-Łojasiewicz (KL) inequality with an exponent of $\frac{1}{2}$ for convex composite functions. When the KL exponent is $\frac{1}{2}$, KL inequality is known as PL inequality for nonsmooth functions. Li and Pong [11] studied the calculus of the exponent of KL inequality and proved the linear convergence for iterates of the PGM when the PL inequality holds. Recently, Garrigos et al. [10] provided an explicit linear convergence rate of the PGM under the PL inequality. Zhang and Zhang [18] used a proximal map-based residual to define the generalized PL inequality for convex composite functions, and showed an explicit linear convergence rate of the PGM under this inequality.
@@
Drori and Teboulle [8] first proposed the performance estimation problem (PEP) for the worst-case analysis of first-order methods, which is based on using a semidefinite programming (SDP) relaxation. Subsequently, PEP has been widely used for analyzing the worst-case convergence rates of various first-order methods, such as the gradient methods or its accelerated versions [8, 6], the PGM [16], the alternating direction method of multipliers [17], and the difference-of-convex algorithm [1]. While the original SDP relaxation in [8] is based on a discrete relaxation of the PEP,Taylor et. al. [15, 14] proposed an exact reformulation of the PEP by deriving the sufficient and necessary interpolation conditions for the function class $\mathcal{F}_{\mu,L}$ with $\mu\geq 0$ and $L\in(\mu,\infty]$. For a composite function with a strongly convex $f$, Taylor et al. [16] showed the exact worst-case linear convergence rates of the PGM in terms of objective function accuracy, distance to optimality, and residual gradient norm.
@@
The main contribution of this paper is applying performance estimation to analyze the convergence of the PGM for composite functions under two classes of PL inequalities: the PL inequality, the variant of PL inequality defined by the proximal map-based residual, which we call RPL for simplicity. When the PL inequality holds, we provide the first explicit linear convergence rate for nonconvex composite functions, and improve the existing convergence results in [10] for convex composite functions. When the RPL inequality holds, we also provide the first explicit linear convergence rate for nonconvex composite function, and find a better bound than that given in [18] for convex composite functions. Furthermore, for the two classes of PL inequalities, we deduce the “optimal” constant step size based on the convergence rates obtained above. We compared in Table 1 the explicit worst-case convergence rates we obtained with the existing results in the literature.
@@
$f(x)\in\mathcal{F}_{0,L}$, $h(x)\in\mathcal{F}_{0,+\infty}$ PL $\max\{\frac{1}{2\gamma\mu+1},\frac{(L\gamma-1)^{2}}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu}\}$ $\frac{1}{1+\gamma\mu(2-\gamma L)}$ [10]
@@
$f(x)\in\mathcal{F}_{-L,L}$, $h(x)\in\mathcal{F}_{0,+\infty}$ PL $\max\{\frac{(L\gamma+1)^{2}}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu},\frac{(L\gamma-1)^{2}}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu}\}$ /
@@
The rest of the paper is organized as follows. In Sect. 2, we provide preliminary definitions of PL inequalities and some interpolation conditions used in the proof. In Section 3, we present explicit linear convergence rates of the PGM and derive its “optimal” step size under the PL inequality. Subsequently, in Section 4, we extend our analysis to the RPL inequality setting and also establish new linear convergence results. We end the paper with some conclusions in Sect. 5.
@@ 2 Preliminaries
We consider the convergence for the PGM under the following two settings, where the first consider nonconvex smooth function $f$, and the latter considers convex smooth $f$.
@@
The function $f(x)\in\mathcal{F}_{-L,L}$ and $h(x)\in\mathcal{F}_{0,+\infty}$.
@@
The function $f(x)\in\mathcal{F}_{0,L}$ and $h(x)\in\mathcal{F}_{0,+\infty}$.
@@
$\frac{1}{2\mu}||\nabla f(x)||^{2}\geq(f(x)-f_{*}),\quad\forall x\in\mathbb{R}^{d}.$ (2)
The PL inequality implies that each stationary point is a global minimum.
Let $\partial F(x)$ denote the Clarke subdifferential of a local Lipschitz function $F(x)$ We say the function $F$ satisfies the PL inequality if there exists $\mu>0$ such that
$\frac{1}{2\mu}d(0,\partial F(x))^{2}\geq F(x)-F_{*},\quad\forall x\in\mathbb{R}^{d},$ (3)
where $\partial F(x)=\nabla f(x)+\partial h(x)$ because $f$ is smooth [4, Proposition 2.3.3]. Note that the PL inequality here is a special case of the KL inequality ([3]) with exponent $\frac{1}{2}$.
For composite function $F(x)$, Zhang and Zhang [18] used proximal map-based residual to define the RPL inequality:
@@
Our analysis heavily depends on the PEP technique, which can be summarised as the following procedure: (1) The abstract PEP problem can be discretized as the finite-dimensional optimization problem, where interpolation conditions should be used if available. (2) By introducing the Gram matrix, the finite-dimensional optimization problem is relaxed as a convex SDP problem, which is solvable via various off-the-shelf solvers. (3) Optimal dual solutions are first obtained numerically, and then the analytical expressions are guessed and verified analytically. (4) Finally, the worst-case convergence rate is obtained by using nonnegative linear combinations of constraints in the discretized formulation of the PEP problem.
The following facts play a key role in our analysis. Consider a set of data triples $\{(x_{i},g_{i},f_{i})\}_{i\in I}$ and define the index set $I=\{1,\ldots,N,*\}$ for the iterations. Recall the following interpolation conditions for two classes of functions.
@@
[14, Theorem 3.10] The set $\{(x_{i};g_{i};f_{i})\}_{i\in I}$ is $\mathcal{F}_{-L,L}$-interpolable if and only if the following inequality holds $\forall i,j\in I$:
@@
[15, Corollary 1] The set $\{(x_{i};g_{i};f_{i})\}_{i\in I}$ is $\mathcal{F}_{0,L}$-interpolable if and only if the following inequality holds $\forall i,j\in I$:
@@
[15, Theorem 1] The set $\{(x_{i};s_{i};h_{i})\}_{i\in I}$ is $\mathcal{F}_{0,\infty}$-interpolable if and only if the following set of conditions holds $\forall i,j\in I$:
@@ 3.1 Nonconvex Composite Case
In this section, we consider the performance of the PGM under Condition 1 and assume the PL inequality (3) holds. In the following of this paper, we set the performance criterion as the objective function accuracy $F(x_{2})-F_{*}$. In this setting, the PEP for the PGM can be written as follows:
@@
where $\Delta\geq 0$ denotes the difference between the optimal function value $F_{*}$ and the value of $F$ at the starting point. In addition, $F$ and $x_{1}$ are decision variables. This is an infinite-dimensional optimization problem with infinite number of constraints, and consequently intractable in general. In Section 2, Facts 1 and 3 provide necessary and sufficient conditions for the interpolation of $L$-smooth functions and closed, proper and convex functions. We can discretize problem (8) as a finite-dimensional optimization problem by using the following (in)equalities and interpolation conditions (5) and (7) in Facts 1 and 3.
First, following [16], the iteration $x_{k+1}=\mathrm{prox}_{\gamma h}\left(x_{k}-\gamma\nabla f(x_{k})\right)$ can be rewritten using necessary and sufficient optimality conditions on the definition of the proximal operator:
$x_{k+1}=x_{k}-\gamma(g_{k}+s_{k+1})$ (9)
for some $s_{k+1}\in\partial h(x_{k+1})$, and $g_{k}=\nabla f(x_{k})$. Note that $g_{*}+s_{*}=0$.
@@
Suppose that Condition 1 holds, $F$ satisfy the PL inequality (3) on $X=\{x:F(x)\leq F(x_{1})\}$, and $x_{2}$ is generated by Algorithm 1. Then the following holds.
@@
1. First consider the case of $\gamma\in\left(0,\frac{\sqrt{3}}{L}\right]$. Let $\alpha_{1}=\frac{L^{2}\gamma^{2}+3L\gamma+2}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu},\alpha_{2}=\frac{L\gamma+1}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu},$ $\beta_{1}=\rho=\frac{(L\gamma+1)^{2}}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu}$, and $\lambda_{1}=\frac{L\gamma^{2}\mu+2\gamma\mu}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu}.$ It is obvious that $\alpha_{1},\alpha_{2},\beta_{1},\lambda_{1}$, and $\rho$ are nonnegative. We obtain that
@@
where $A_{i,j}$ is defined in (5), $C_{1,2}$ is defined in (7), $D_{2}$ is defined in (10), and the last inequality holds since $\gamma\in(0,\frac{\sqrt{3}}{L}]$. Because the terms after the nonnegative parameters $\alpha_{1},\alpha_{2},\beta_{1},\lambda_{1}$ are nonpositive, we have $F(x_{2})-F_{*}\leq(\alpha_{1}-\alpha_{2})(F(x_{1})-F_{*}).$ We then obtain the desired result by substituting the values of $\alpha_{1}$ and $\alpha_{2}$.
2. For the case of $\gamma\in\left(\frac{\sqrt{3}}{L},\frac{2}{L}\right)$, let $\alpha_{1}=\frac{L\gamma-1}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu},\alpha_{2}=\frac{-L^{2}\gamma^{2}+3L\gamma-2}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu},$ $\beta_{1}=\rho=\frac{(L\gamma-1)^{2}}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu},\lambda_{1}=\frac{-L\gamma^{2}\mu+2\gamma\mu}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu}.$ It is obvious that $\alpha_{1},\alpha_{2},\beta_{1},\lambda_{1}$, and $\rho$ are nonnegative. By multiplying the inequalities by their respective factors and then summing them, we obtain that
@@
where $A_{i,j}$ is defined in (5), $C_{1,2}$ is defined in (7), $D_{2}$ is defined in (10), and the last inequality holds since $\gamma\in(\frac{\sqrt{3}}{L},\frac{2}{L})$. The remainder of the proof follows a similar approach to the first case. ∎
We remark that the parameters $\alpha_{1},\alpha_{2},\beta_{1},\lambda_{1}$, and $\rho$ are obtained from the SDP relaxation of the PEP problem, which is derived based on investigating the dual optimal solutions from numerous numerical experiments. Note that the explicit dual feasible solution and convergence rate in Theorem 1 match those of the SDP relaxation from numerical experiments.
@@
Existing literature has studied the linear convergence properties of gradient methods for nonconvex and nonsmooth functions under the PL inequality (3), e.g, Li and Pong [11] showed that if $F(x)$ satisfies the PL inequality (3), the iterative sequence $\{x_{k}\}$ converges locally linearly to a stationary point of $F(x)$. However, to the best of our knowledge, there is no explicit rate of linear convergence in existing literature. In Theorem 1, we provide the first explicit linear convergence rate of the PGM on objective function values for convex composition minimization.

### v2

@@ Abstract
We study the linear convergence rates of the proximal gradient method for composite functions satisfying two classes of Polyak-Łojasiewicz (PL) inequality: the PL inequality, the variant of PL inequality defined by the proximal map-based residual. Using the performance estimation problem, we either provide new explicit linear convergence rates or improve existing complexity bounds for minimizing composite functions under the two classes of PL inequality. Finally, we illustrate numerically the effects of our theoretical results.
@@ 1 Introduction
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is a function with Lipschitz continuous gradient, and $h:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is a closed, proper and convex function. In addition, we assume that the optimal set $X_{*}(F):=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}F(x)$ is nonempty. We denote the class of closed and proper convex functions by $\mathcal{F}_{0,\infty}(\mathbb{R}^{d})$, the class of $L$-smooth convex functions by $\mathcal{F}_{0,L}(\mathbb{R}^{d})$, and the class of $L$-smooth functions by $\mathcal{F}_{-L,L}(\mathbb{R}^{d})$. For simplicity, we say that $F$ is a nonconvex composite function if $f\in\mathcal{F}_{-L,L}(\mathbb{R}^{d})$ and a convex composite function if $f\in\mathcal{F}_{0,L}(\mathbb{R}^{d})$.
@@
2:for $i=1,\ldots,N$ do
3:     $x_{i+1}=\mathrm{prox}_{\gamma h}\left(x_{i}-\gamma\nabla f(x_{i})\right)$
4:end for
@@
It is well-known that for composite functions with strongly convex $f$, the PGM achieves a linear convergence rate [2]. Relaxing the strong convexity condition while retaining a linear convergence rate of the PGM has been extensively studied in the literature. The PL inequality was initially introduced by Polyak [17], which is used to show linear convergence of gradient descent. Bolte et al. [4] derived a linear convergence rate of the PGM under the Kurdyka-Łojasiewicz (KL) inequality with an exponent of $\frac{1}{2}$ for convex composite functions. When the KL exponent is $\frac{1}{2}$, the KL inequality is known as PL inequality for nonsmooth functions. Li and Pong [13] studied the calculus of the exponent of KL inequality and proved the linear convergence for the iterates of the PGM when the PL inequality holds. Recently, Garrigos et al. [11] provided an explicit linear convergence rate of the PGM under the PL inequality. Zhang and Zhang [23] used a proximal map-based residual to define the generalized PL inequality for convex composite functions, and showed an explicit linear convergence rate of the PGM under this inequality.
@@
Drori and Teboulle [9] first proposed the performance estimation problem (PEP) for the worst-case analysis of first-order methods, which is based on using a semidefinite programming (SDP) relaxation. Subsequently, PEP has been widely used for analyzing the worst-case convergence rates of various first-order methods, such as the gradient methods or their accelerated versions [9, 7], the PGM [20], the alternating direction method of multipliers [22], and the difference-of-convex algorithm [1]. While the original SDP relaxation in [9] is based on a discrete relaxation of the PEP, Taylor et. al. [19, 18] proposed an exact reformulation of the PEP by deriving the sufficient and necessary interpolation conditions for the function class $\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$ with $\mu\geq 0$ and $L\in(\mu,+\infty]$. For a composite function with a strongly convex $f$, Taylor et al. [20] showed the exact worst-case linear convergence rates of the PGM in terms of objective function accuracy, distance to optimality, and residual gradient norm.
@@
The main contribution of this paper is applying performance estimation to analyze the convergence of the PGM for composite functions under two classes of PL inequalities: the PL inequality, the variant of PL inequality defined by the proximal map-based residual, which we call RPL for simplicity. When the PL inequality holds, we provide the first explicit linear convergence rate for nonconvex composite functions, and improve the existing convergence results in [11] for convex composite functions. When the RPL inequality holds, we also provide the first explicit linear convergence rate for nonconvex composite functions, and find a better bound than that given in [23] for convex composite functions. Furthermore, for the two classes of PL inequalities, we deduce the “optimal” constant step size based on the convergence rates obtained above. We compare in Table 1 the explicit worst-case convergence rates we obtained with the existing results in the literature.
@@
$f\in\mathcal{F}_{0,L}(\mathbb{R}^{d})$, $h\in\mathcal{F}_{0,\infty}(\mathbb{R}^{d})$ PL $\max\{\frac{1}{2\gamma\mu+1},\frac{(L\gamma-1)^{2}}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu}\}$ $\frac{1}{1+\gamma\mu(2-\gamma L)}$ [11]
@@
$f\in\mathcal{F}_{-L,L}(\mathbb{R}^{d})$, $h\in\mathcal{F}_{0,\infty}(\mathbb{R}^{d})$ PL $\max\{\frac{(L\gamma+1)^{2}}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu},\frac{(L\gamma-1)^{2}}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu}\}$ /
@@
The rest of the paper is organized as follows. In Sect. 2, we provide preliminary definitions of PL inequalities and some interpolation conditions used in the proof. In Sect. 3, we present explicit linear convergence rates of the PGM and derive its “optimal” step size under the PL inequality. Subsequently, in Sect. 4, we extend our analysis to the RPL inequality setting and also establish new linear convergence results. In Sect. 5, we verify the performance of the “optimal” step size in practical problems through numerical experiments. We end the paper with some conclusions in Sect. 6.
@@ 2 Preliminaries
We consider the convergence for the PGM under the following two settings, where the first considers nonconvex smooth function $f$, and the latter considers convex smooth $f$.
@@
The function $f\in\mathcal{F}_{-L,L}(\mathbb{R}^{d})$ and $h\in\mathcal{F}_{0,\infty}(\mathbb{R}^{d})$.
@@
The function $f\in\mathcal{F}_{0,L}(\mathbb{R}^{d})$ and $h\in\mathcal{F}_{0,\infty}(\mathbb{R}^{d})$.
@@
$||\nabla f(x)||^{2}\geq 2\mu(f(x)-f_{*}),\quad\forall x\in\mathbb{R}^{d},$ (2)
where $f_{*}$ denotes the global optimal function value. The PL inequality implies that each stationary point is a global minimum.
Let $\partial F(x)$ denote the Clarke subdifferential of a local Lipschitz function $F$ at $x$ [5]. We say the function $F=f+h$ satisfies the PL inequality if there exists $\mu>0$ such that
$d(0,\partial F(x))^{2}\geq{2\mu}(F(x)-F_{*}),\quad\forall x\in\mathbb{R}^{d},$ (3)
where $\partial F(x)=\nabla f(x)+\partial h(x)$ because $f$ is smooth [5, Proposition 2.3.3]. Note that the PL inequality here is a special case of the KL inequality with exponent $\frac{1}{2}$ [4].
For composite function $F$, Li and Li [14] used proximal map-based residual to define the RPL inequality:
@@
Lemma 1.
[10, Theorem 3.5] Consider the composite function $F=f+h$. For any $x\in\mathbb{R}^{d}$, it holds that $\|\mathcal{G}_{\gamma}(x)\|\leq d(0,\partial F(x)).$
By Lemma 1, it is clear that the RPL inequality (4) implies the PL inequality (3), since
$F(x)-F_{*}\leq\frac{1}{2\mu}\|\mathcal{G}_{\gamma}(x)\|^{2}\leq\frac{1}{2\mu}d(0,\partial F(x))^{2}.$
However, it appears challenging to derive the converse implication from the PL inequality to the RPL inequality.
Our analysis heavily depends on the PEP technique, which can be summarised as the following procedure: (1) The abstract PEP problem can be discretized as a finite-dimensional optimization problem, where interpolation conditions should be used if available. (2) By introducing the Gram matrix, the finite-dimensional optimization problem is relaxed as a convex SDP problem, which is solvable via various off-the-shelf solvers. (3) Optimal dual solutions are first obtained numerically, and then the analytical expressions are guessed and verified analytically. (4) Finally, the worst-case convergence rate is obtained by using nonnegative linear combinations of constraints in the discretized formulation of the PEP problem.
We denote first-order oracle at a given point $x$ for function $F$ by $\mathcal{O}_{F}(x)=\{x,F(x),\tilde{\nabla}F(x)\}$, where ~$\tilde{\nabla}F(x)\in\partial F(x)$ is a Clarke subgradient of $F$ at $x$ (or gradient if $F$ is differentiable at $x$). We consider a method $\mathcal{M}$ that executes $N$ steps of iterations $\{x_{i}\}_{1\leq i\leq N}$ starting from the initial iterate $x_{1}$, where the $i+1$-th iteration generated by $\mathcal{M}$ can be written as $x_{i+1}=\mathcal{M}_{i+1}(x_{1},\mathcal{O}_{F}(x_{1}),\ldots,\mathcal{O}_{F}(x_{i}))$. Let $I=\{1,\ldots,N,*\}$ be an index set containing an optimal solution $x_{*}$ of the problem. At each iterate $x_{i}$, we denote the output of the oracle by $\{x_{i},g_{i},f_{i}\}$, where $f_{i}=F(x_{i})$ and $g_{i}\in\partial F(x_{i})$. Based on the above concepts, we introduce the definition of $\mathcal{F}$-interpolation.
Definition 1.
[19, Definition 2] Let $I=\{1,\ldots,N,*\}$ be an index set, and consider the set of triples $\{(x_{i},g_{i},f_{i})\}_{i\in I}$, where $x_{i}\in\mathbb{R}^{d}$, $g_{i}\in\mathbb{R}^{d}$ and $f_{i}\in\mathbb{R}$, $\forall i\in I$. The set of triples $\{(x_{i},g_{i},f_{i})\}_{i\in I}$ is $\mathcal{F}$-interpolable if and only if there exists a function $F\in\mathcal{F}$ such that for all $i\in I$, we have both $f_{i}=F(x_{i})$ and $g_{i}\in\partial F(x_{i})$.
Recall the following interpolation conditions, which play a key role in our analysis.
@@
[18, Theorem 3.10] The set $\{(x_{i};g_{i};f_{i})\}_{i\in I}$ is $\mathcal{F}_{-L,L}$-interpolable if and only if the following inequality holds $\forall i,j\in I$:
@@
[19, Corollary 1] The set $\{(x_{i};g_{i};f_{i})\}_{i\in I}$ is $\mathcal{F}_{0,L}$-interpolable if and only if the following inequality holds $\forall i,j\in I$:
@@
[19, Theorem 1] The set $\{(x_{i};s_{i};h_{i})\}_{i\in I}$ is $\mathcal{F}_{0,\infty}$-interpolable if and only if the following set of conditions holds $\forall i,j\in I$:
@@ 3.1 Nonconvex Composite Case
We now consider the performance of the PGM under Condition 1 and assume the PL inequality (3) holds. In the following of this paper, we set the performance criterion as the objective function accuracy $F(x_{2})-F_{*}$. In this setting, the PEP for the PGM can be written as follows:
@@
where $X\subseteq\mathbb{R}^{d}$, $\Delta\geq 0$ denotes the difference between the optimal function value $F_{*}$ and the value of $F$ at the starting point. In addition, $F$ and $x_{1}$ are decision variables. This is an infinite-dimensional optimization problem with an infinite number of constraints, and consequently intractable in general. In Section 2, Facts 1 and 3 provide necessary and sufficient conditions for the interpolation of $L$-smooth functions and closed, proper and convex functions. We can discretize problem (8) as a finite-dimensional optimization problem by using the following (in)equalities and interpolation conditions (5) and (7) in Facts 1 and 3.
First, following [20], the iteration $x_{i+1}=\mathrm{prox}_{\gamma h}\left(x_{i}-\gamma\nabla f(x_{i})\right)$ can be rewritten using necessary and sufficient optimality conditions on the definition of the proximal operator:
$x_{i+1}=x_{i}-\gamma(g_{i}+s_{i+1})$ (9)
for some $s_{i+1}\in\partial h(x_{i+1})$, and $g_{i}=\nabla f(x_{i})$. Note that $g_{*}+s_{*}=0$.
@@
Suppose that Condition 1 holds, $F$ satisfies the PL inequality (3) on $X=\{x:F(x)\leq F(x_{1})\}$, and $x_{2}$ is generated by Algorithm 1. Then the following holds.
@@
1. First consider the case of $\gamma\in\left(0,\frac{\sqrt{3}}{L}\right]$. Let $\alpha_{1}=\frac{L^{2}\gamma^{2}+3L\gamma+2}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu},\alpha_{2}=\frac{L\gamma+1}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu},$ $\beta_{1}=\frac{(L\gamma+1)^{2}}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu}$, and $\lambda_{1}=\frac{L\gamma^{2}\mu+2\gamma\mu}{(L\gamma+1)^{2}+L\gamma^{2}\mu+2\gamma\mu}.$ It is obvious that $\alpha_{1},\alpha_{2},\beta_{1}$, and $\lambda_{1}$ are nonnegative. We obtain that
@@
where $A_{1,2}$ and $A_{2,1}$ are defined in (5), $C_{1,2}$ is defined in (7), $D_{2}$ is defined in (10), and the last inequality holds since $\gamma\in(0,\frac{\sqrt{3}}{L}]$. Because the terms after the nonnegative parameters $\alpha_{1},\alpha_{2},\beta_{1},\lambda_{1}$ are nonpositive, we have $F(x_{2})-F_{*}\leq(\alpha_{1}-\alpha_{2})(F(x_{1})-F_{*}).$ We then obtain the desired result by substituting the values of $\alpha_{1}$ and $\alpha_{2}$.
2. For the case of $\gamma\in\left(\frac{\sqrt{3}}{L},\frac{2}{L}\right)$, let $\alpha_{1}=\frac{L\gamma-1}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu},\alpha_{2}=\frac{-L^{2}\gamma^{2}+3L\gamma-2}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu},$ $\beta_{1}=\frac{(L\gamma-1)^{2}}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu},\lambda_{1}=\frac{-L\gamma^{2}\mu+2\gamma\mu}{(L\gamma-1)^{2}-L\gamma^{2}\mu+2\gamma\mu}.$ It is obvious that $\alpha_{1},\alpha_{2},\beta_{1}$, and $\lambda_{1}$ are nonnegative. By multiplying the inequalities by their respective factors and then summing them, we obtain that
@@
where $A_{1,2}$ and $A_{2,1}$ are defined in (5), $C_{1,2}$ is defined in (7), $D_{2}$ is defined in (10), and the last inequality holds since $\gamma\in(\frac{\sqrt{3}}{L},\frac{2}{L})$. The remainder of the proof follows a similar approach to the first case. ∎
We remark that the parameters $\alpha_{1},\alpha_{2},\beta_{1}$, and $\lambda_{1}$ are obtained from the SDP relaxation of the PEP problem, which is derived based on investigating the dual optimal solutions from numerous numerical experiments. Note that the explicit dual feasible solution and convergence rate in Theorem 1 match those of the SDP relaxation from numerical experiments.
@@
Existing literature has studied the linear convergence properties of gradient methods for nonconvex and nonsmooth functions under the PL inequality (3), e.g, Li and Pong [13] showed that if $F$ satisfies the PL inequality (3), the iterative sequence $\{x_{k}\}$ converges locally linearly to a stationary point of $F$. However, to the best of our knowledge, there is no explicit rate of linear convergence in existing literature. In Theorem 1, we provide the first explicit linear convergence rate of the PGM on objective function values for convex composition minimization.